OpenStudy (anonymous):

an investment of $10,000 in 1980 Hershey stock was worth $536,000 in 1998. Assuming continuous growth, what was the annual growth rate?

OpenStudy (anonymous):

\[\frac{536,000-10,000}{1998-1980}=\frac{526,000}{18}\approx29,222.22\]Thus the annual growth rate is: \[\frac{\frac{526,000}{18}}{18}=\frac{526,000}{18}*\frac{1}{18}=\frac{526,000}{18²}\approx1623.46\] Not sure though.